Suppose that all investors expect that interest rates for the 4 years will be as follows:
...

Suppose that all investors expect that interest rates for the 4 years will be as follows:


Year 0 (today): 4 %

Year 1: 5 %

Year 2: 6 %

Year 3: 7 %

What is the price of a 2-year maturity bond with a 10% coupon rate paid annually? (Par value = $1,000)
Answer #1

Given that,

rate for current year r1 = 4%

Rate for next year r2 = 5%

For a 2-year 10% coupon bond, with face value of $1000,

price = C/(1+r1) + (FV+C)/((1+r1)*(1+r2))

where C is annual coupon = 10% of 1000 = $100

So, price = 100/1.04 + (1000+100)/(1.04*1.05) = $1103.48

So, price of the bond is $1103.48